# The Entropic Skin: Spatial Entanglement from the QCD Confinement Boundary

### v1

@@ II.1 The Ground State Spinor Profile
$\psi(\mathbf{r})=N\begin{pmatrix}ij_{0}(Er)\chi_{s}\\ -j_{1}(Er)(\bm{\sigma}\cdot\hat{\mathbf{r}})\chi_{s}\end{pmatrix}$, (1)
@@ II.2 The Local Reduced Density Matrix
To quantify the radial dependence of spin-position entanglement, we construct the local reduced density matrix for the spin subsystem, $\rho_{S}(r)$. We trace out the angular coordinates $(\theta,\phi)$ at a fixed radius $r$:
$\rho_{S}(r)=\frac{1}{\mathcal{P}(r)}\int d\Omega\,\text{Tr}_{\text{spatial}}\left(|\psi\rangle\langle\psi|\right),$ (2)
where $\mathcal{P}(r)$ is the local probability density required to normalize $\text{Tr}(\rho_{S})=1$.
Due to the rotational symmetry of the $s$-wave ground state, the off-diagonal elements vanish upon integration. For an initial spin-up state ($\chi_{\uparrow}$), the density matrix becomes diagonal:
@@
Here, $P_{\downarrow}(r)$ represents the probability that the spin has “flipped” relative to the $z$-axis. This flip is not a dynamic transition but a geometric requirement of the lower component spinor structure.
@@ II.4 The Two Limits
Spin-Flip Probability ($P_{\downarrow}$) 0 $1/3$
@@ III Discussion: The Fate of the Entropic Skin
Physically, this surface entropy can be understood as a coordinate-space realization of the vacuum entanglement observed in high-energy experiments. By imposing a sharp boundary condition at $r=R$, the Bag Model effectively partitions the QCD vacuum, severing the correlations between the interior valence region and the exterior. The “Entropic Skin” of $\approx 0.918$ bits represents the irreducible uncertainty injected into the quark state by this vacuum partitioning. It acts as the geometric projection of the severed vacuum correlations.
The existence of this “Entropic Skin” raises a fundamental question about the restoration of Chiral Symmetry. In the static Bag Model, the chirality mixing at the boundary is hard-wired. However, in Chiral Bag Models (like the Cloudy Bag [Thomas1984, Owa2022]), the axial current is conserved by the emission of a pion field at the surface.

### v2

@@ II.1 The Ground State Spinor Profile
$\psi(\mathbf{r})=N\begin{pmatrix}j_{0}(Er)\chi_{s}\\ -ij_{1}(Er)(\bm{\sigma}\cdot\hat{\mathbf{r}})\chi_{s}\end{pmatrix}$, (1)
@@ II.2 The Local Reduced Density Matrix
To quantify the radial dependence of spin-position entanglement, we construct the local reduced density matrix for the spin subsystem, $\rho_{S}(r)$. Decomposing the Dirac spinor into upper ($\phi$) and lower ($\chi$) 2-spinors, $\psi=(\phi,\chi)^{T}$, we trace out the angular degrees of freedom and sum over the Dirac components:
$\rho_{S}(r)=\frac{1}{\mathcal{N}(r)}\int d\Omega\left[\phi(\mathbf{r})\phi^{\dagger}(\mathbf{r})+\chi(\mathbf{r})\chi^{\dagger}(\mathbf{r})\right].$ (2)
Here, $\mathcal{N}(r)$ is the radial probability density, which serves as the local normalization factor to ensure $\text{Tr}(\rho_{S})=1$:
$\mathcal{N}(r)=\int d\Omega\,\psi^{\dagger}(\mathbf{r})\psi(\mathbf{r})\propto j_{0}^{2}(xr)+j_{1}^{2}(xr).$ (3)
Due to the rotational symmetry of the $s$-wave ground state, the off-diagonal elements of the density matrix vanish upon integration. For an initial spin-up state ($\chi_{\uparrow}$), $\rho_{S}(r)$ becomes diagonal:
@@
Here, $P_{\downarrow}(r)$ represents the probability that the spin has flipped relative to the quantization axis due to the orbital angular momentum contribution of the lower component. This mixing corresponds to the transfer of angular momentum from intrinsic spin to the orbital angular momentum of the lower component, a mechanism central to discussions of the proton spin budget [Myhrer2010].
@@ II.4 The Two Limits
Spin-Flip Probability ($Ptheoreticalproposals_{\downarrow}$) 0 $1/3$
@@ III Discussion: The Fate of the Entropic Skin
Physically, this surface entropy can be understood as a coordinate-space realization of the vacuum entanglement observed in high-energy experiments, consistent with recent proposals of an entropic horizon derived from the QCD trace anomaly [Mamo:2025hur]. By imposing a sharp boundary condition at $r=R$, the Bag Model effectively partitions the QCD vacuum, severing the correlations between the interior valence region and the exterior. The “Entropic Skin” of $\approx 0.918$ bits represents the irreducible uncertainty injected into the quark state by this vacuum partitioning. It acts as the geometric projection of the severed vacuum correlations.
The existence of this “Entropic Skin” raises a fundamental question about the restoration of Chiral Symmetry. In the static Bag Model, the chirality mixing at the boundary is hard-wired. However, in Chiral Bag Models (like the Cloudy Bag [PhysRevD.22.2838, PhysRevD.24.216, Thomas1984, Owa2022]), the axial current is conserved by the emission of a pion field at the surface.